OpenStudy (anonymous):

The mean, mode, median, range of 5 11 9 5 25

OpenStudy (anonymous):

mean =( 5+11+9+5+25 ) / 5 = 55/ 5 = 11 mode = 5 (5 occurs most) median = 9

OpenStudy (anonymous):

range is largest - smalles 25 - 5 = 20

OpenStudy (anonymous):



OpenStudy (anonymous):

Mean =11 , mode=5 , median =9, range =20